# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v4

@@ 5.2 Data Collection
Effective risk assessment relies on both internal and external data sources to capture a comprehensive picture of potential risks. Internal data from ERP systems, procurement records, and inventory management systems provides insights into operational processes. External data, such as market indicators, social media sentiment, and climate reports, are critical for Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. ERP systems, supplier systems, and logistics systems, alongside external sources like social media and weather forecasts. Such data covers both structured (e.g., numerical data) and unstructured formats (e.g., text or multimedia) to enhance risk detection. Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. The author used Twitter streaming API as a primary source for data extraction. Kosasih et al., (2022) utilized external platforms like Marklines for automotive SC data and Achilles Information Ltd for energy sector insights. Pan and Miao, (2023) collected data through surveys from industry practitioners to capture qualitative and contextual insights about SCR.Including such data enhances the model’s ability to preemptively detect risks like demand fluctuations or reputational issues, underscoring the value of a multimodal data approach.
@@ 5.4 Model Selection
Practitioners have a range of model choices, each suited to different data characteristics and objectives in SCRA. Classical models like Support Vector Machine (SVM), Decision Tree (DT), and Logistic Regression(LR) are ideal for well-structured, low-dimensional datasets, where the data relationships are more straightforward and computational efficiency is prioritized. DL models, such as Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, are valuable for handling complex, high-dimensional data, including time-series demand data, where capturing temporal dependencies is essential. Hybrid models, which combine conventional machine learning (CML) and DL approaches, balance interpretability and accuracy, catering to scenarios where insights into model decisions are as important as predictive performance. Ensemble methods like Random Forest(RF) and XGBoost aggregate outputs from multiple models to enhance predictive accuracy and robustness. Additional research emphasizes that model selection should consider factors such as data availability, computational resources, and the necessity for model interpretability. Highly interpretable models are particularly important in regulated industries, such as finance, where transparency is critical. Hybrid and ensemble models, increasingly popular in recent literature, often effectively compromise predictive power and interpretability, making them suitable for complex SC risk scenarios.
@@ 5.8 Preventive Measure
Chen et al., (2024) GA-BPNN, Logistic Regression High accuracy in predicting MSME default and credit ratings (Accuracy: 92% for default prediction, 86% for credit rating).
@@ 6.2 AI Models Used in SCRA
The application of AI in SCRA has seen significant growth, driven by the need for robust, predictive models to manage risk in complex and dynamic supply networks. To effectively analyze the methodologies employed in the reviewed literature, we classified the AI models into five main groups: CML Models, DL Models, Hybrid Models, Ensemble Methods, and Advanced Ensemble Methods. Each category is defined based on its distinct methodological characteristics, level of complexity, application scope, and suitability for specific SC risk scenarios, as shown in Figure 15.
@@
6.2.1 Conventional ML (CML) Models
Support Vector Machine (SVM): SVM is a supervised ML algorithm effective in high-dimensional spaces. It finds a hyperplane to separate classes in the feature space, often using kernel functions for non-linear separation. SVM performs well with limited data, making it useful for SCs with restricted historical data. By maximizing the margin, SVMs are less prone to overfitting, especially in smaller datasets. However, SVM can be computationally expensive for large datasets, which can be limiting in SCs with extensive data. Its results are also difficult to interpret, especially with complex kernel functions, which can be challenging for explainable predictions in SC risk. Bassiouni et al., (2023) found that the SVM classifier achieved an accuracy of 84.56% in predicting SC risks. Fu et al., (2022) found that the SVM model achieved a classification accuracy of 95.31% in assessing credit risk in Internet-based SC financing, outperforming the BPNN but slightly underperforming the BP-GA model. SVM is ideal for high-stakes, high-dimensional risk scenarios, like financial risk assessment with limited historical data. The data type is structured, often with limited samples. Input variables include supplier attributes, transaction history, and financial indicators, with the target variable being binary risk classification (e.g., high-risk vs. low-risk suppliers).
Decision Tree (DT): DTs are a tree-based algorithm that splits data into branches based on feature values, resulting in a flowchart-like model that is easy to interpret. DTs offer a simple structure that is easy to interpret, making it valuable in SC risk scenarios that require transparency. They are also relatively fast to train and execute, which makes them suitable for real-time or near-real-time assessments. However, DT models are prone to overfitting, especially with noisy data, and may require pruning or ensemble methods to mitigate this. They also struggle with capturing complex, non-linear relationships compared to models like neural networks. Rao and Li, (2022) found that the DT model achieved an accuracy of 77.2% in predicting SC financing risks. However, while highly interpretable, Baryannis et al., 2019a found that DT learning achieved lower performance predicting SC risks than SVM. The DT model demonstrated a trade-off between interpretability and performance, with SVM providing superior predictive accuracy, particularly in binary classification tasks for late deliveries. DT models are adequate for straightforward, interpretable assessments of SC risk, especially with structured and smaller datasets. The data type is structured with categorical or binary risk indicators. Input variables include financial ratios, transaction amounts, and supplier location, with the target variable being risk level classification (e.g., high, medium, low).
Logistic Regression (LR): LR is a linear model that estimates probabilities for categorical outcomes, commonly used for binary classification in risk assessment. LR offers a straightforward, easy-to-interpret linear approach that is beneficial for transparent decision-making in SC risk. It also requires less data and computational power, making it suitable for smaller datasets. However, LR is unsuited for complex, non-linear relationships, which can reduce accuracy in intricate SC risk scenarios. It also assumes feature independence, which might not hold in interconnected SC networks. Li, (2022) found that the LR model achieved a high prediction accuracy of 92.8% in forecasting investment risks in SCM, demonstrating excellent applicability and potential for broader use in SCM risk prediction. While Yang et al., (2021) found that the Logistic Regression model achieved a prediction accuracy of 83.1% in evaluating the credit risk of SMEs in SC finance, the lasso logistic model significantly outperformed it with an accuracy of 96.5% and a lower Type II error rate. LR is ideal for binary risk scenarios and cases where model transparency is essential. The data type is structured, typically tabular data. Input variables include supplier payment history, financial indicators, and supply volume, with the target variable being binary risk classification (e.g., default vs. no default).
6.2.2 Deep Learning (DL) Models
DL models, like neural networks, use multiple layers to model complex, non-linear relationships, making them practical for complex SC risk predictions. DL models are ideal for intricate, non-linear relationships in SC data, such as demand forecasting under uncertain conditions. They generally achieve high accuracy in well-established scenarios, particularly with large datasets. However, the “black-box” nature of DL models can make them difficult to interpret, which may limit adoption in risk-sensitive contexts. They also require large datasets and computational power, hindering some SC applications. Pan and Miao, (2023) found that the BPNN model demonstrated excellent performance in SCRA, achieving a maximum relative error of only 0.0021%, significantly outperforming the AHP model, which had a maximum relative error of 57.41%. This indicates the BPNN’s superior ability to evaluate risk accurately and reliably in SC systems. DL is suitable for large-scale, complex SCRAs with high data availability and less need for interpretability. The data type is large, structured, or unstructured data. Input variables include time series data on SC metrics, inventory levels, and demand trends, with the target variable being the probability of disruption or a risk score.
6.2.3 Hybrid Models (CML + DL Models)
Hybrid models combine CML and DL techniques to leverage the strengths of each for more robust SCRA. Hybrid models often outperform single models by combining strengths, such as using DL for feature extraction and CM for interpretability. They are versatile across various scenarios by tailoring components to specific data characteristics and risk types. However, hybrid models can be challenging to implement and tune, requiring expertise in multiple algorithms. Additionally, combining models can increase computational costs, which may be challenging for real-time applications. Fu et al., (2022) found that the hybrid BPNN and GA model outperformed both the standalone BPNN and SVM models in predicting credit risk for internet-based SC financing. The BPNN-SVM hybrid achieved the highest classification accuracy of 97.19%, significantly improving prediction performance compared to the BPNN (82.33%) and SVM (95.31%) models. Hybrid models are well-suited for applications requiring both high accuracy and moderate interpretability, especially when handling complex, multi-source data. The data type is mixed, including structured and semi-structured data. Input variables include supplier reliability, lead times, transactional volumes, and customer demand, with the target variable being multi-class risk categorization (e.g., high, medium, low).
6.2.4 Ensemble Methods
Ensemble methods like Bagging, Boosting, and Stacking combine multiple models to improve accuracy and robustness in SC risk prediction. Ensemble methods reduce variance and bias by aggregating multiple models, which makes them suitable for high-stakes risk assessment. They are versatile, as they can work with various algorithms, allowing for tailored approaches depending on the data and risk scenario. However, ensemble methods can require substantial computational resources, particularly for Boosting and Stacking. While they boost accuracy, they can decrease interpretability, making it difficult to understand individual model contributions. Liu and Huang, (2020) utilized the EN-AdaPSVM model, integrating AdaBoost with SVM and enhanced by noise elimination and adaptive mutation particle swarm optimization, achieving high classification accuracy and robustness in SC finance risk assessment. Luo et al., (2022) highlighted that while their AdaBoost-DPSO-SVM model showed significant performance improvements, it faced challenges such as higher error rates in negative sample recognition, suggesting potential issues with bias and misclassification. Ensemble methods are ideal for high-accuracy, high-stakes SC risk scenarios, especially when model interpretability is not the primary concern. The data type is structured, tabular data, often large datasets. Input variables include financial history, supplier performance scores, and geographic risk indicators, with the target variable being binary or multi-class risk levels (e.g., high-risk, low-risk, moderate risk).
6.2.5 Advanced Ensemble Methods
XGBoost: XGBoost is a powerful gradient-boosting algorithm known for its speed and accuracy, especially in structured data applications. XGBoost often outperforms traditional algorithms in structured data tasks, providing robust risk predictions. It also handles missing data well, which is useful in SC datasets that may be incomplete. However, XGBoost models can be complex, reducing interpretability and making tuning more challenging. Liu et al., (2022) developed a credit risk assessment model for the construction industry’s online SC finance using a hybrid approach combining XGBoost, SMOTENC, and RF. The study aimed to address the gaps in existing credit assessment systems, which often rely on traditional models with low accuracy and poor interpretability. Using data from 71 upstream enterprises between 2018 and 2020, the model achieved an overall accuracy rate of 91.67% on test samples. XGBoost is effective for structured, high-dimensional datasets in SC risk, particularly where interpretability is secondary to accuracy. The data type is large, structured data, often tabular. Input variables include financial indicators, transactional history, and geographic data, with the target variable being risk classification or score.
Random Forest (RF): RF is an ensemble method based on DTs, known for reducing variance and improving predictive accuracy. RF retains interpretability compared to more complex models and effectively handles high-dimensional data while being less prone to overfitting. However, RF can become slow with large datasets, particularly when building large forests. High-depth trees can also lead to overfitting if not carefully tuned. Wang et al., (2022) utilized a cost-sensitive Random Forest (CSL-RF) model to predict SMEs’ credit risk in SC finance, incorporating 40 diverse predictors and employing an imbalance sampling strategy for high accuracy. Athaudage et al., (2022) applied RF regression to forecast crude oil prices during disease outbreaks by integrating disease data with oil market variables, demonstrating reliable performance under complex conditions. Both studies highlight the effectiveness of RF models in handling varied data sets for robust forecasting. RF is effective when interpretability is moderately important and high accuracy and robustness are needed. The data type is structured, often large-scale. Input variables include supplier attributes, financial ratios, and location-based risk factors, with the target variable being multi-class risk level (e.g., high, medium, low).
@@ 7 Practical and Managerial Implications
The findings of this article will play a significant role in managerial strategies designed to enhance SCRM. First, it stresses that ML and AI are important tools that will give rise to a new phase of SCRA. Managers are encouraged to try advanced ML models, such as RF, XGBoost, and blends of these. They stimulate considerable risk assessment precision. These models are likely to enhance the accuracy of risk evaluation enormously and, thus, the effectiveness of their risk management strategies.
@@
In addition, the study highlights the importance of flexibility and adaptability in an era after COVID-19, suggesting that companies need new SC strategies. Managers must now place contingency plans at the forefront, focusing on those not readily broken by any future disruptions. It is only by doing routine work thoroughly that one can concentrate on something special without fear or disturbance from the outside world, using people who have already developed a variety of fresh paths, so avoiding crisis after crisis with all that effort having been wasted.
@@
The findings also highlight the importance of ensemble methods – RF and XGBoost – and hybrid methodologies in risk assessment. It is possible that such models could improve the accuracy of risk prediction and may be of interest to managers striving to strengthen their SCRM systems. The importance of time series prediction models, such as ARIMA and LSTM, cannot be overstated. Companies are advised to create tools predicting potential risks based on historical performance to approach the task of SC disruption mitigation with much-needed foresight.
@@
XAI and ML models are extremely important. Models must produce precise and human-readable outputs as they assist decision-makers in understanding the process of risk assessment and even its findings. Furthermore, the study indicates the significance of real-time data mining and using non-traditional sources like social media, e.g., Twitter, as critical components of full-scale risk identification. Managers should be careful and up-to-date in understanding real-time information.
@@
Blockchain represents an effective instrument for increasing transparency and traceability in the SC, thereby making risk monitoring mechanisms more efficient. Organizations exploring blockchain solutions should determine the possibilities of addressing and using the risk to boost SCR. SMEs require principal strategies in risk evaluation in terms of CR since they have distinct characteristics and risks in terms of operations. Conversely, implementing strategies based on SMEs’ needs and constraints is required. Moreover, federated learning should be implemented when data privacy issues arise. This method addresses the risk but protects confidential information, permitting partners to conduct joint risk forecasting. Compliance with security and data privacy must be performed almost immediately.
@@
Several hybrid risk assessment methodologies that combine various ML models and techniques bring the necessary comprehensive view of SCRA, offering managers the chance to look into such approaches to develop risk management strategies. Additionally, comparing benchmarks of different ML models is critical to understanding more information about the best ones for particular risk scenarios. The overall message of the study is outlining the need for continuous performance evaluations and improvements. Risk assessment strategies should be able to evolve alongside other circumstances. Other assumptions refer to integrating ML-based risk assessment into already existing SC workflows and decision-making patterns. It is necessary to determine how such technologies can be easily implemented. Employee training in data analysis, ML, and AI is critical. Managers should invest in skill development to empower their SC professionals to harness these technologies effectively. Collaboration and knowledge sharing within organizations and across industries are encouraged to collectively build robust risk management strategies. Such partnerships have the potential to enhance risk resilience significantly.
@@
Lastly, understanding and adhering to data protection and privacy regulations is critical when deploying AI and ML for risk assessment. Managers should ensure that their risk management practices align with the legal requirements to avoid compliance issues. The study offers a comprehensive roadmap for organizations aiming to fortify their SCRM strategies in an ever-evolving landscape.

### v5

@@ 5.2 Data Collection
Effective risk assessment relies on internal and external data sources to capture a comprehensive picture of potential risks. Internal data from ERP systems, procurement records, and inventory management systems provides insights into operational processes. External data, such as market indicators, social media sentiment, and climate reports, are critical for Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. ERP systems, supplier systems, and logistics systems, alongside external sources like social media and weather forecasts. Such data covers both structured (e.g., numerical data) and unstructured formats (e.g., text or multimedia) to enhance risk detection. Deiva Ganesh and Kalpana, (2022) used real-time data from social media, that is, Twitter, as the input source. The author used Twitter streaming API as a primary source for data extraction. Kosasih et al., (2022) utilized external platforms like Marklines for automotive SC data and Achilles Information Ltd for energy sector insights. Pan and Miao, (2023) collected data through surveys from industry practitioners to capture qualitative and contextual insights about SCR.Including such data enhances the model’s ability to preemptively detect risks like demand fluctuations or reputational issues, underscoring the value of a multimodal data approach.
@@ 5.4 Model Selection
Practitioners have a range of model choices, each suited to different data characteristics and objectives in SCRA. Classical models like Support Vector Machine (SVM), Decision Tree (DT), and Logistic Regression (LR) are ideal for well-structured, low-dimensional datasets, where the data relationships are more straightforward and computational efficiency is prioritized. DL models, such as Convolutional Neural Networks (CNNs) and Long Short-Term Memory (LSTM) networks, are valuable for handling complex, high-dimensional data, including time-series demand data, where capturing temporal dependencies is essential. Hybrid models, which combine conventional ML and DL approaches, balance interpretability and accuracy, catering to scenarios where insights into model decisions are as important as predictive performance. Ensemble methods like Random Forest (RF) and XGBoost aggregate outputs from multiple models to enhance predictive accuracy and robustness. Additional research emphasizes that model selection should consider data availability, computational resources, and the necessity for model interpretability. Highly interpretable models are particularly important in regulated industries, such as finance, where transparency is critical. Hybrid and ensemble models, increasingly popular in recent literature, often effectively compromise predictive power and interpretability, making them suitable for complex SC risk scenarios.
@@ 5.8 Preventive Measure
Chen et al., (2024) GA-BPNN, LR High accuracy in predicting MSME default and credit ratings (Accuracy: 92% for default prediction, 86% for credit rating).
Zhang et al., (2025) Domain-Adaptation-based Multistage Ensemble Learning (DAMEL) DEMEL outperformed traditional models in SME credit risk evaluation, addressing data scarcity, distribution discrepancies, and class imbalance (AUC: 0.7652).
Zhou and Wang, (2025) XGBoost, RF, CNN, Least Squares SVM (LSSVM) XGBoost outperformed all models with accuracy: 91.9%, precision: 92.8%, recall: 94.7%, F1-score: 93.7%. SHAP analysis showed asset-liability ratio, quick ratio, and ESG ratings as key risk indicators.
Mojdehi et al., (2025) BallMapper + Graph Neural Network (BM-GNN) BM-GNN effectively captures financial indicators and network-based features, improving SME credit risk assessment in SC Finance.
@@ 6.2 AI Models Used in SCRA
A wide range of ML and AI techniques have been applied to SC risk prediction, reflecting the increasingly data-driven nature of modern risk management. These techniques can be broadly classified into five clusters: Classical ML (CML) models, DL models, Hybrid CML+DL approaches, Ensemble methods, and Emerging Techniques such as Federated Learning, Reinforcement Learning, and Large Language Models (LLMs). Each cluster offers distinct advantages, limitations, and application contexts, with the literature providing diverse empirical evidence supporting their use. Each category is defined based on its distinct methodological characteristics, level of complexity, application scope, and suitability for specific SC risk scenarios, as shown in Figure 14.
@@
6.2.1 CML Models
CML models remain popular in SC risk prediction due to their simplicity, interpretability, and computational efficiency. Models such as Support Vector Machines (SVM), Logistic Regression (LR), Decision Trees (DT), and Bayesian Networks (BN) are particularly effective when working with structured, tabular data where relationships between variables are relatively well understood. For example, SVM and its variants have been used successfully in several studies for credit risk assessment, supplier default prediction, and operational risk evaluation (Zhu et al.,, 2016; Zhang et al.,, 2015; Zhao and Li,, 2022; Li and Fu,, 2022; Zhang et al.,, 2021; Liu and Huang,, 2020; Zhang et al.,, 2022; Fu et al.,, 2022; Liu et al.,, 2022; Xia et al.,, 2023; Lei et al.,, 2023; Gao et al.,, 2022; Yao et al.,, 2022). These models, often enhanced by feature selection or optimization algorithms, show strong classification accuracy, particularly when combined with techniques such as genetic algorithms (GA) and principal component analysis (PCA). Logistic regression remains a favored baseline due to its transparency and ease of interpretation, applied in risk assessment settings where explainability is required, such as in Li, (2022); Rao and Li, (2022); Chen et al., (2024). Decision Trees (DT), employed in studies like Baryannis et al., 2019a ; Rao and Li, (2022), provide interpretable decision rules but often suffer from overfitting unless combined into ensembles. Meanwhile, Bayesian Networks (BN) offer a probabilistic, causal perspective on risk propagation, particularly in disruption analysis settings, as demonstrated by Hosseini and Barker, (2016). Lasso regression models, which excel in feature sparsity, were applied in Yang et al., (2021) to isolate key predictors of SC risk. ML-based linear regression algorithm (ML-LRA) proposed by Wei, (2022) reduces supplier credit risk. Though highly interpretable, these models often struggle with capturing non-linear relationships and require significant domain expertise to manually engineer relevant features, limiting their scalability in high-dimensional or unstructured data environments.
Advantages
  • 1.
    Simple implementation and training.
  • 2.
    Interpretability supports regulatory compliance.
  • 3.
    Suitable for smaller datasets.
Disadvantages
  • 1.
    Limited ability to capture complex, non-linear patterns.
  • 2.
    May require extensive feature engineering.
Typical Applications
  • 1.
    Early-stage risk identification.
  • 2.
    Interpretable models for regulatory audits.
  • 3.
    Situations with limited historical data.
6.2.2 DL Models
In contrast, DL models have demonstrated superior performance in capturing complex, non-linear patterns in SC risk prediction, particularly when large, high-dimensional, or unstructured data. LSTM networks, known for their strength in sequence modeling, have been widely applied for time series forecasting in SC finance and disruption prediction (Chen et al.,, 2021; Zhang et al.,, 2023; Janjua et al.,, 2023; Gao et al.,, 2022). CNN has also been adopted, particularly when analyzing spatially structured data or feature maps, as seen in Yin et al., (2022); Zhou and Wang, (2025); Zhang et al., (2023). Multimodal DL (Zhang,, 2022; Bassiouni et al.,, 2023; Bodendorf et al.,, 2023; Pan and Miao,, 2023) models have emerged to integrate diverse data types, such as financial indicators, operational metrics, and external news sources, to enhance predictive performance, with DeepRisk showing strong results in Zhang, (2022). More recently, tensorization-based neural symbolic models (Katsaliaki et al.,, 2022) have been proposed, blending symbolic reasoning with DL for relational risk prediction. While DL models demonstrate state-of-the-art accuracy, their application in SCs is hampered by high computational costs, data-hungry training processes, and their opaque, black-box nature, which hinders regulatory acceptance in some industries.
Advantages
  • 1.
    Superior performance with complex, high-dimensional data.
  • 2.
    Effective at handling temporal and spatial dependencies.
  • 3.
    Potential to combine multimodal data.
Disadvantages
  • 1.
    High computational cost.
  • 2.
    Poor interpretability without dedicated explainability tools.
  • 3.
    Requires large labeled datasets
Typical Applications
  • 1.
    High-frequency SC data (sensor data, IoT).
  • 2.
    Multimodal data integration (text, time series, images).
  • 3.
    Long-term risk forecasting.
6.2.3 Hybrid Models (CML + DL)
Hybrid approaches combining classical ML and DL techniques aim to strike a balance between interpretability and predictive accuracy, particularly in scenarios where data characteristics vary across dimensions or stages of the risk prediction process. Several studies combined genetic algorithms (GA) with Backpropagation Neural Networks (BPNN), resulting in enhanced convergence rates and prediction precision (Wu et al., 2022a, ; Fu et al.,, 2022; Nezamoddini et al.,, 2020; Duan et al.,, 2021; Chen et al.,, 2024). In Jianying et al., (2021), PSO was also used to optimize BPNN parameters (Dang et al.,, 2022; Han and Zhang,, 2021), outperforming single models. Hybrid model chains, such as the XGBoost-SMOTENC-RF pipeline proposed in Liu et al., (2022), improve performance under class imbalance scenarios common in SC risk datasets. A particularly novel hybrid, TD-LSTM combined with deep neural networks (DNN), demonstrated strong forecasting accuracy in SC financial data in Zhang et al., (2023). By combining domain-specific feature engineering with data-driven learning, these hybrid approaches address non-linearity while preserving interpretability, though they introduce higher complexity and tuning costs.
Advantages
  • 1.
    Balance between interpretability and predictive power.
  • 2.
    Flexibility to handle different data types and distributions.
  • 3.
    Often more robust to overfitting.
Disadvantages
  • 1.
    Increased model complexity.
  • 2.
    Longer training times.
  • 3.
    Risk of suboptimal tuning in heterogeneous frameworks.
Typical Applications
  • 1.
    Medium-sized datasets with mixed data types.
  • 2.
    Cases requiring both prediction accuracy and interpretability.
  • 3.
    Data with substantial noise and missing values.
6.2.4 Ensemble Models (Bagging, Boosting, Stacking)
Ensemble methods, particularly RF (Athaudage et al.,, 2022; Wang et al.,, 2022, 2021; Xia et al.,, 2023; Liu et al.,, 2022; Zhu et al.,, 2019), boosting algorithms (XGBoost, AdaBoost) (Liu et al.,, 2022; Zhou and Wang,, 2025; Zhang et al.,, 2025), and multi-boosting approaches consistently outperform individual SC risk prediction models by aggregating predictions from multiple weak learners. Studies such as Zhu et al., (2019, 2017); Liu and Huang, (2020); Belhadi et al., 2021b demonstrated that multi-boosting ensembles effectively handle data noise and feature redundancy, improving accuracy in risk classification tasks. Cost-sensitive Random Forest (CSL-RF) models further enhance robustness under unbalanced risk categories, as shown in Wang et al., (2022). Novel approaches like the Domain-Adaptation-based Multistage Ensemble Learning (DAMEL) framework, proposed in Zhang et al., (2025), directly address data scarcity, distribution shifts, and class imbalance, particularly relevant for SME credit risk prediction. While ensembles excel in predictive accuracy, their major trade-off is reduced interpretability, especially as ensemble size increases.
Advantages
  • 1.
    Enhanced predictive performance.
  • 2.
    More robust to overfitting.
  • 3.
    Can combine strengths of multiple algorithms.
Disadvantages
  • 1.
    Computationally expensive.
  • 2.
    Poor interpretability.
  • 3.
    Difficult to tune optimally.
Typical Applications
  • 1.
    Large, noisy datasets.
  • 2.
    Situations where prediction accuracy is paramount.
  • 3.
    Multi-source data environments.
6.2.5 Emerging and Hybrid Techniques
Several cutting-edge techniques have emerged, demonstrating strong potential for real-time risk monitoring and collaborative prediction across SC partners. Federated learning (FL) approaches, such as those in Zheng et al., (2023), enable decentralized risk modeling across firms without exposing sensitive data, a crucial advantage in privacy-sensitive SC networks. Reinforcement Learning (RL), employed in Aboutorab et al., (2022), optimizes adaptive intervention strategies, balancing risk mitigation costs and performance outcomes. Graph Neural Networks (GNN), such as the BM-GNN framework introduced in Mojdehi et al., (2025), excel at capturing relational dependencies between suppliers, enriching risk propagation modeling. Lastly, natural language processing (NLP) and LLMs, including GPT-4 and BERT, have been applied for text-based risk identification, extracting risk signals from news, social media, and operational reports (Wu,, 2024; Xu et al.,, 2024; Deiva Ganesh and Kalpana,, 2022).
Advantages
  • 1.
    Enables collaborative modeling across multiple organizations.
  • 2.
    Preserves data privacy.
  • 3.
    Hybrid symbolic-ML interpretability.
Disadvantages
  • 1.
    Requires careful coordination across participants.
  • 2.
    Vulnerable to communication bottlenecks.
  • 3.
    Model updates may suffer from drift across sites.
Typical Applications
  • 1.
    Collaborative supplier monitoring.
  • 2.
    Risk prediction across supply networks.
  • 3.
    Situations where data privacy regulations restrict data sharing.
@@ 7 Practical and Managerial Implications
The findings of this article will play a significant role in managerial strategies designed to enhance SCRM. First, it stresses that ML and AI are important tools that will give rise to a new phase of SCRA. Managers are encouraged to try advanced ML models, such as RF, XGBoost, and blends. They stimulate considerable risk assessment precision. These models are likely to enhance the accuracy of risk evaluation enormously and, thus, the effectiveness of their risk management strategies.
@@
In addition, the study highlights the importance of flexibility and adaptability in an era after COVID-19, suggesting that companies need new SC strategies (Sheng et al.,, 2021). Managers must now place contingency plans at the forefront, focusing on those not readily broken by future disruptions. It is only by doing routine work thoroughly that one can concentrate on something special without fear or disturbance from the outside world, using people who have already developed a variety of fresh paths, so avoiding crisis after crisis with all that effort having been wasted.
@@
The findings also highlight the importance of ensemble methods like RF and XGBoost in risk assessment (Zhou and Wang,, 2025). It is possible that such models could improve the accuracy of risk prediction and may be of interest to managers striving to strengthen their SCRM systems. The importance of time series prediction models, such as ARIMA and LSTM, cannot be overstated. Companies are advised to create tools predicting potential risks based on historical performance to approach the task of SC disruption mitigation with much-needed foresight.
@@
ML models must produce precise and human-readable outputs as they assist decision-makers in understanding the process of risk assessment and even its findings. Furthermore, the study indicates the significance of real-time data mining and using non-traditional sources like social media, e.g., Twitter, as critical components of full-scale risk identification (Deiva Ganesh and Kalpana,, 2022). Managers should be careful and up-to-date in understanding real-time information.
@@
Blockchain, combined with ML models, represents a powerful tool for enhancing transparency, traceability, and risk monitoring in the SC. Blockchain ensures secure and immutable data sharing, while ML models enable predictive risk assessment and anomaly detection, making risk management more proactive and efficient (Dang et al.,, 2022). Organizations exploring blockchain solutions should assess how integrating ML-driven risk forecasting can enhance SC resilience and optimize risk mitigation strategies. SMEs require principal strategies in risk evaluation regarding credit risks since they have distinct characteristics and risks in terms of operations. Conversely, implementing strategies based on SMEs’ needs and constraints is required. Moreover, federated learning should be implemented when data privacy issues arise. This method addresses the risk but protects confidential information, permitting partners to conduct joint risk forecasting. Compliance with security and data privacy must be performed almost immediately.
@@
Several hybrid risk assessment methodologies that combine various ML models and techniques bring the necessary comprehensive view of SCRA (Mojdehi et al.,, 2025). These approaches enable managers to explore advanced risk management strategies, enhancing their ability to effectively identify, assess, and mitigate risks. Additionally, comparing benchmarks of different ML models is critical to understanding more information about the best ones for particular risk scenarios. The overall message of the study is outlining the need for continuous performance evaluations and improvements. Risk assessment strategies should be able to evolve alongside other circumstances. Other assumptions refer to integrating ML-based risk assessment into already existing SC workflows and decision-making patterns. It is necessary to determine how such technologies can be easily implemented. Employee training in data analysis, ML, and AI is critical. Managers should invest in skill development to empower their SC professionals to harness these technologies effectively. Collaboration and knowledge sharing within organizations and across industries are encouraged to collectively build robust risk management strategies. Such partnerships have the potential to enhance risk resilience significantly.
@@
Lastly, understanding and adhering to data protection and privacy regulations is critical when deploying AI and ML for risk assessment. Managers should ensure that their risk management practices align with the legal requirements to avoid compliance issues.